NASA Hubble Space Telescope Daily Report #4441
NASA Hubble Space Telescope Daily Report #4441
http://www.spaceref.com/news/viewsr.nl.html?pid=25336
"SIGNIFICANT EVENTS:NICMOS Recovery Flash Report: NICMOS recovery
procedure was successfully completed at 246/15:03 UTC. This procedure
enabled NICMOS Buffer Box 2 and transitioned NICMOS up to Operate mode
and then back down to SAAOPER. The mode transition was performed in
order to refresh the Buffer Box (BB) telemetry. The current temperatures
for BB1, 2 and 3 are -1.94, -7.76 and -2.42 respectively.
NICMOS is now configured to intercept science at the next available
opportunity. The nominal plan is to intercept science at the beginning of
next week's SMS (Day 253)."
